# Mass diffusion and bending in dynamic wetting by phase-field and sharp interface models

### v2

@@ V.1 Sheared drop
Figure 2: Correlation between the radius of curvature $R$ and (41). (a) Each simulation corresponds to a unique set of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers while maintaining a constant capillary number ($\operatorname{Ca}$). The geometrical factor $a$ remains fixed at 2.5, and a fitted line demonstrates a clear one-to-one relationship between these two parameters. (b) Steady-state solutions of phase-field (PF) simulations for cases highlighted in red in the left panel.
In the first test case, we study the steady shape of a nanoscopic droplet confined between two moving walls similarly to Lacis et al. [19, 20]. Initially, we use the phase-field (PF) model to solve the system in order to validate the correlation between the radius of curvature and the flow parameters, as expressed in (41). We choose a constant capillary number $\operatorname{Ca}=0.0212$, which is below the critical capillary number, while systematically varying the Peclet and Cahn numbers. The geometrical factor $a$ is kept fixed at a value of 2.5. For each simulation run, we compute the radius of curvature $R$ once the system attains a steady state, using the mean curvature at the contact line. The results, illustrated in Figure 2, demonstrate a remarkable agreement between $R$ and (41), establishing a clear one-to-one relationship between these two parameters. This preliminary analysis reinforces our choice of the geometrical factor $a$. In Section V.2, we will further substantiate the appropriateness of this choice in the context of a spreading droplet.
We now test our toy model on this setup. Using the relations described in Section IV, we compare full PF simulations – in the sense that the full domain is solved – with VOF simulations using the toy model approximation. In the toy model, the bottom boundary is now a computational boundary that is located a distance $+\delta$ away from the wall. The domain comprised between the wall $y=0$ and the computational boundary $y=\delta$ is not simulated. There, we apply the dynamic angle relation (43), where the apparent angle $\theta_{\mathrm{a}}$ is now a dynamic contact angle $\theta_{d}$
@@
with $U_{w}$ the wall velocity. Figure 3 shows the resulting steady state interfaces for the left side of a sheared droplet, for a range of $\operatorname{Pe}$ and $\operatorname{Cn}$. The full PF simulations are compared to the corresponding VOF results, using equations (44) and (45) as boundary conditions, with good agreement. A more discerning test is shown in Figure 4 where the local angle that the interface makes with the horizontal is shown as a function of vertical position, for the same six cases as in Figure 3. The details of the profile shapes are near identical in all cases.
@@
Figure 4: Comparison between the phase-field angle along the left interface (black points) and the volume-of-fluid angle (red line) for six distinct sets of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers.
@@ V.2 Spreading drop
In this second test case, we study a simple drop spreading over a no-slip substrate. We impose an equilibrium angle, denoted as $\theta_{\mathrm{e}}$, which governs the final shape of the drop at equilibrium. Contrary to the previous case, we need to relax the impermeability condition by allowing a controlled mass flux through the computational boundary. This introduced vertical sinking velocity, denoted as $v_{\mathrm{s}}$, is linked to the net mass change, being either a loss (in the case of $\theta_{\mathrm{e}}<\theta_{0}$) or a gain (in the case of $\theta_{\mathrm{e}}>\theta_{0}$) of mass. The boundary condition governing the normal component of velocity at the wall is now
$u_{y}|{y=\delta}=v_{\mathrm{s}}.$ (46)
@@
The initial radius of the drop is $R_{0}=0.5$ with an initial contact angle $\theta_{0}=90\degree$. The equilibrium angle is set to $\theta_{\mathrm{e}}=70\degree$. The dimensionless parameters governing the flow are $\operatorname{Ca}=0.0212$, $\operatorname{Re}=3.978$, $\operatorname{Cn}=0.01$ and $\operatorname{Pe}=1$. Figure 5 shows the influence of varying the parameters $a$ and $b$ for this transient case. Each panel shows a close up of the left contact line region, with PF results for three different times, and the corresponding VOF results, using a range of values for a and b. The best match is found for $a=3$ and $b=0.2170$ confirming the values chosen in the previous case.
@@
Figure 6 shows the diameter of the drop as a function of time for the PF and the VOF simulations. In order to compare the displacement, the position of the contact line in the VOF setup is projected back to the wall $y=0$ using the extracted apparent angle located at $y=\delta+\Delta$, in order to avoid the difference in imposed-extracted angles $\Delta\theta_{\mathrm{num}}$ at $y=\delta$. From the comparison between both models, we can observe that the displacement curves superimpose almost perfectly. The final equilibrium position of the contact line is the same in both cases, implying that the sinking velocity $v_{\mathrm{sink}}$ allowed us to accurately represent the mass flux through the computational boundary. In the transient state, however, we see a slight difference between times $0.1$ and $0.3$ where the VOF simulation overshoots the solution. This behavior is similar to ones observed previously in the super-slip model and is related to the effect of the slip.
Figure 7: Phase-field and toy model volume-of-fluid interfaces at times t = 0, 0.05, 0.1, 0.15, 0.2 and 0.4. The dotted black line corresponds to the location of the computational boundary in the toy model.
Figure 8: Zoom around the contact line on the phase-field and toy model volume-of-fluid interfaces at times t = 0, 0.05, 0.1, 0.15, 0.2 and 0.4. The dotted black line corresponds to the location of the computational boundary in the toy model.
Figure 7 shows the interfaces for both models at times 0, 0.05, 0.1, 0.15, 0.2 and 0.4, and Figure 8 . The blue line corresponds to the location of the computational boundary. In the VOF simulation, the domain is cut at $y=\delta$ and anything below it is not simulated. The superposition of the PF and VOF interfaces show good agreement. In Figure 8, we show the same interfaces zoomed at the contact line. The first observation is the appearance of a discontinuity in the last VOF fragment. Again, this showcases the numerical difference $\Delta\theta_{\mathrm{num}}$, where the actual angle is always different from the one imposed through height functions in the ghost layer. Nevertheless, the VOF interfaces show a relatively good agreement with respect to the PF ones.
@@ VI Conclusion
The phase field model for dynamic wetting simulations has the advantage that it can be explicitly connected to a particular physical situation, i.e. that of a mixture of two near immiscible fluids, and thus in principle only depends on physically meaningful parameters. Computationally it is however quite expensive, and it is not suited for large scale CFD applications. On the other hand, the volume-of-fluid method which is the method of choice for such simulations, depends on ad hoc or numerical parameters for contact line motion.
In this paper we have attempted to bridge this gap by making a local analysis of the phase field equations near the contact line, establishing the connection between a local strong curvature and the contact line speed. VOF simulations are generally successful at capturing the interface in the bulk of the liquid, and it is only in the region near the contact line that the special features of PF are important. Using this local analysis we have parameterized the PF results, and put them in the form of a boundary condition that can be used in VOF simulations of moving contact lines. We thus propose that this will allow VOF simulations to give results that are equivalent in many respects to the results of a full PF simulation, but at the cost of a VOF simulation.
@@ Appendix A Sink velocity
In the toy model, a computational boundary is introduced at a distance $+\delta$ from the wall where we apply a Navier boundary condition. When considering a spreading drop, the impermeability condition ($u_{y}|_{y=\delta}=0$) needs to be relaxed in order to take into account the mass flux through that computational boundary. To that end, we introduce a sink velocity $v_{\mathrm{sink}}$, such that
@@
[…]
@@
[…]
@@
[…]

### v4

@@ V.1 Sheared drop
Figure 2: Correlation between the radius of curvature $R$ and (41). (a) Each simulation corresponds to a unique set of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers while maintaining a constant capillary number ($\operatorname{Ca}$). The toy model parameters are set to $a=2.5$ and $b=0.2170$, and the fitted line demonstrates a clear one-to-one relationship between these two quantities. (b) Steady-state solutions of phase-field (PF) simulations for cases highlighted in red in the left panel.
In the first test case, we study the steady shape of a nanoscopic droplet confined between two moving walls similarly to Lacis et al. [lacis_steady_2020, lacis2022]. Initially, we use the phase-field model to solve the system in order to validate the correlation between the radius of curvature and the flow parameters, as expressed in (41). We choose a constant capillary number $\operatorname{Ca}=0.0212$, which is below the critical capillary number, while systematically varying the Peclet and Cahn numbers. The equilibrium angle is set to $\theta_{e}=90\degree$ for both the advancing and receding contact lines, and the geometrical factor $a=2.5$ is kept fixed. For each simulation run, we compute the radius of curvature $R$ once the system reaches steady state, using the mean curvature at the contact line. The results, illustrated in Fig. 2, demonstrate a remarkable agreement between $R$ and (41), establishing a clear one-to-one relationship between these two quantities. In this preliminary analysis, the parameter $b=0.2170$ is fixed (see Appendix B for details). In Section V.2, we will further substantiate the appropriateness of this choice in the context of a spreading droplet.
We now test our toy model on this setup. Using the relations described in Section IV, we compare full PF simulations – in the sense that the full domain is solved – with VOF simulations using the toy model approximation. In the toy model, the bottom boundary is now a computational boundary that is located a distance $\delta$ away from the wall. The domain comprised between the wall $y=0$ and the computational boundary $y=\delta$ is not simulated. There, we apply the dynamic angle relation (43), where the apparent angle $\theta_{\mathrm{a}}$ is now a dynamic contact angle $\theta_{d}$
@@
with $U_{w}$ the wall velocity. Fig. 3 shows the resulting steady state interfaces for the left side of a sheared droplet, for a range of $\operatorname{Pe}$ and $\operatorname{Cn}$. The full PF simulations are compared to the corresponding VOF results, using equations (44) and (45) as boundary conditions, with good agreement. A more discerning test is shown in Fig. 4 where the local angle that the interface makes with the horizontal is shown as a function of vertical position, for the same six cases as in Fig. 3. The details of the profile shapes are near identical in all cases.
@@
Figure 4: Comparison between the phase-field angle along the left interface (black points) and the volume-of-fluid angle (red symbols) for six distinct sets of Peclet ($\operatorname{Pe}$) and Cahn ($\operatorname{Cn}$) numbers.
@@ V.2 Spreading drop
In this second test case, we study a simple drop spreading over a no-slip substrate. We impose an equilibrium angle which governs the final shape of the drop at equilibrium. Contrary to the previous case, we need to relax the impermeability condition by allowing a controlled mass flux through the computational boundary. This introduced vertical sinking velocity, denoted as $v_{\mathrm{s}}$, is linked to the net mass change, being either a loss (in the case of $\theta_{\mathrm{e}}<\theta_{0}$) or a gain (in the case of $\theta_{\mathrm{e}}>\theta_{0}$) of mass, with $\theta_{0}$ the initial angle of the drop. The boundary condition governing the normal component of velocity at the wall is now
$\left.u_{y}\right|_{y=\delta}=v_{\mathrm{s}}.$ (46)
@@
The initial radius of the drop is $R_{0}=0.5$ with an initial contact angle $\theta_{0}=90\degree$. The equilibrium angle is set to $\theta_{\mathrm{e}}=70\degree$. The dimensionless parameters governing the flow are $\operatorname{Ca}=0.0212$, $\operatorname{Re}=3.978$, $\operatorname{Cn}=0.01$ and $\operatorname{Pe}=1$. Fig. 5 shows the influence of varying the parameters $a$ and $b$ for this transient case. Each panel shows a close up of the left contact line region, with PF results for three different times, and the corresponding VOF results, using a range of values for $a$ and $b$. The best match is found for $a=3$ and $b=0.2170$ confirming the values chosen in the previous case. Note that the difference in $a$ between both cases is attributed to the different equilibrium contact angle.
@@
Fig. 6 shows the diameter of the drop as a function of time for the PF and the VOF simulations. In order to compare the displacement, the position of the contact line in the VOF setup is projected back to the wall $y=0$ using the extracted apparent angle located at $y=\delta+\Delta$, in order to avoid the numerical difference between imposed and extracted angles $\Delta\theta_{\mathrm{num}}$ at $y=\delta$. From the comparison between both models, we can observe that the displacement curves superimpose almost perfectly. The final equilibrium position of the contact line is the same in both cases, implying that the sinking velocity $v_{\mathrm{sink}}$ allowed us to accurately represent the mass flux through the computational boundary. In the transient state, however, we see a slight difference between times $0.1$ and $0.3$ where the VOF simulation overshoots the solution. This behavior is similar to ones observed previously in slip models [fullana_simulation_2022].
Figure 7: Phase-field and toy model volume-of-fluid interfaces at times t=0, 0.05, 0.1, 0.15, 0.2, 0.4t=0,\;0.05,\;0.1,\;0.15,\;0.2,\;0.4. The blue line corresponds to the location of the computational boundary in the toy model.
Figure 8: Zoom around the contact line of the phase-field and toy model volume-of-fluid interfaces at times t=0, 0.05, 0.1, 0.15, 0.2, 0.4t=0,\;0.05,\;0.1,\;0.15,\;0.2,\;0.4. The blue line corresponds to the location of the computational boundary in the toy model.
Fig. 7 shows both the PF and VOF interfaces at t=0, 0.05, 0.1, 0.15, 0.2, 0.4t=0,\;0.05,\;0.1,\;0.15,\;0.2,\;0.4, and Fig. 8 shows a zoom of the same interfaces near the contact line. The solid blue line indicates the VOF computational boundary: as in the previous case, the domain is truncated at $y=\delta$, so the region $0<y<\delta$ is not simulated. The superposition of the PF and VOF interfaces demonstrates overall good agreement in position and shape. A closer look in Fig. 8 reveals a discontinuity in the final VOF segment close to the contact line. This feature manifests the numerical discrepancy $\Delta\theta_{\mathrm{num}}$, the difference between the contact angle actually reconstructed by the VOF procedure and the angle imposed via the height functions in the ghost layer.
@@ VI Conclusion
The phase-field model for dynamic wetting simulations has the advantage that it can be explicitly connected to a particular physical situation, i.e. that of a mixture of two near immiscible fluids, and thus in principle only depends on physically meaningful parameters. Computationally it is however quite expensive, and it is not suited for large scale CFD applications. On the other hand, the volume-of-fluid method which is the method of choice for such simulations, depends on ad hoc or numerical parameters for contact line motion.
In this paper we have attempted to bridge this gap by making a local analysis of the phase-field equations near the contact line, establishing the connection between a local strong curvature and the contact line speed. VOF simulations are generally successful at capturing the interface in the bulk of the liquid, and it is only in the region near the contact line that the special features of the PF are important. Using this local analysis we have parameterized the PF results, and put them in the form of a boundary condition that can be used in VOF simulations of moving contact lines. We thus propose that this will allow VOF simulations to give results that are equivalent in many respects to the results of a full PF simulation, but at the cost of a VOF simulation. We note in Appendix C that the common case of a single component phase-change fluid can be treated in a similar way, and that an equivalent model for the contact angle can be derived, see equation (69).
Data availability
The data that support the findings of this article are openly available [fullana_2025_16927267].
Acknowledgements
S.Z. and T.F. thank the FLOW laboratory of KTH for hosting them and Shervin Bagheri for funding the visits in 2019, which allowed this collaboration to start. S.Z. was funded by the European Research Council (ERC) through the European Union’s Horizon 2020 Research and Innovation Programme (Grant Agreement No. 883849 TRUFLOW).
@@ Appendix A Sink velocity
In the toy model, a computational boundary is introduced at a distance $\delta$ from the wall where we apply a Navier boundary condition. When considering a spreading drop, the impermeability condition ($u_{y}|_{y=\delta}=0$) needs to be relaxed in order to take into account the mass flux through that computational boundary. To that end, we introduce a sink velocity $v_{\mathrm{sink}}$, such that
@@
We validate the method by considering a drop spreading over a substrate for different values of $\delta$. The drop is initialized with a diameter $D=1$ in a $2\times 2$ domain and we apply a constant contact angle $\theta_{\mathrm{e}}=70\degree$. The viscosities, densities and the surface tension are set to 1. The simulations are run until the equilibrium position is reached. In Fig. 9, we show the total mass as a function of time when applying the sink velocity for $\delta=0.2,\>0.05,\>0.02$. The initial mass inside the drop decreases as spreading occurs. In Fig. 10, we compare the final interfaces for three cases: (i) toy model with sink velocity (ii) toy model without sink velocity (iii) full domain. From the results, we can observe the sink velocity allows us to match to the final true equilibrium shape (full domain). The error in final interface is a function of $\delta$. As $\delta$ is increased, the linear extrapolation procedure – that does not take into account the curved interface at the contact line – induces a higher difference in final shapes. In practice, when using the toy model on a real setup, $\delta$ is small compared to the domain size.
@@
Figure 9: Total mass of the drop as a function of time with or without sink velocity for different computational boundaries placed at $\delta$ with $\delta=0.02,\>0.05,\>0.2$.
@@
Figure 10: Validation of the sink velocity imposed at the computational boundary placed at $\delta$ with $\delta=0.02,\>0.05,\>0.2$ denoted by the dotted black lines in the figures. The equilibrium angle is $\theta_{\mathrm{e}}=70\degree$ for the three cases considered.
Appendix B Toy model calculation of parameter $b$
Rescaling the steady version of (23) using a length-scale $\ell$ and a magnitude $\phi\sim\operatorname{Pe}\ell$ according to (28), we get
$\mathbf{u}\cdot\tilde{\nabla}C=\tilde{\nabla}^{2}\tilde{\phi}$, (47)
where $\tilde{\nabla}$$\ell\nabla$ and $\tilde{\phi}=\phi/(\operatorname{Pe}\ell)$. This equation is thus universal in the sense that all dependencies of $\operatorname{Ca},\operatorname{Pe},\operatorname{Cn}$ etc are scaled out and equation 47 could be solved once and for all. The only remaining independent input parameter is the equilibrium contact angle $\theta_{e}$. The correspondingly rescaled version of the momentum equation will introduce also the ratio of the viscosities in the two fluid regions.
In order to arrive at a simple quantitative model, the dimensionless order unity parameter $b$ is introduced in equation (39). This parameter is expected to essentially capture the equivalent of solving eq 47, and we thus expect that it may depend on $\theta_{e}$ and the viscosity ratio, but that it is independent of $\operatorname{Ca},\operatorname{Pe},\operatorname{Cn}$ etc.
As a bold attempt at estimating $b$ the following model equation was analyzed:
$\operatorname{Pe}\delta\dfrac{dC}{dx}=\dfrac{d^{2}}{dx^{2}}(-\dfrac{d^{2}C}{dx^{2}}-\frac{Cn}{R}\dfrac{dC}{dx}+\Psi^{\prime}(C)),$ (48)
Similarly to the treatment in (30)–(37), we consider a perturbation solution around the equilibrium solution in (35). The equation for the first order perturbation becomes:
$\frac{\operatorname{Pe}\delta}{\operatorname{Cn}}\dfrac{dC_{0}}{dx}+\frac{1}{R}\dfrac{d^{3}C_{0}}{dx^{3}}=\dfrac{d^{2}}{dx^{2}}\left(-\dfrac{d^{2}C_{1}}{dx^{2}}+\Psi^{\prime\prime}(C_{0})C_{1}\right),$ (49)
Multiplying both sides by an $F(x)$ such that $\dfrac{dF(x)}{dx}=C_{0}(x)$, and repeatedly integrating the right hand side partially, it is seen that the right hand side vanishes, yielding a solvability condition on the remaining terms:
$\frac{\operatorname{Pe}\delta}{\operatorname{Cn}}\int F\dfrac{d^{2}F}{dx^{2}}dx=\frac{1}{R}\int F\dfrac{d^{4}F}{dx^{4}}dx$ (50)
Rearranging this, using (39), as
$\frac{\operatorname{Pe}\delta}{\operatorname{Cn}/R}=\frac{1}{2b}\frac{1}{2}\frac{2\sqrt{2}}{3}=\dfrac{\int F\dfrac{d^{4}F}{dx^{4}}dx}{\int F\dfrac{d^{2}F}{dx^{2}}dx}$ (51)
The integrals on the right hand side can both be calculated, giving $\int F\dfrac{d^{4}F}{dx^{4}}dx=\dfrac{2\sqrt{2}}{3}$ and $\int F\dfrac{d^{2}F}{dx^{2}}dx=2\sqrt{2}(1-\ln 2)$ .
This yields
$b=\frac{1-\ln 2}{\sqrt{2}}=0.2170$ (52)
It should be emphasized that the derivation of this number is by no means a rigorous proof, and that $b$, as well as the parameter $a$ introduced in (40), are to be adjusted according to comparisons between the model in (41)–(43), and full solutions of the phase field equations, as shown in Fig. 5. As seen there however, the value $b=0.2170$ gives satisfactory results, and it has been used throughout.
Appendix C Single component fluid with phase change.
In this appendix we sketch a similar development for the case of a single component fluid with phase change, i.e. a volatile droplet in its own vapor. The van der Waals equation of state would be a first example.
Instead of (16) and (17), we will need to begin from the full compressible equations, including the energy equation
$\dfrac{\partial\rho}{\partial t}+(\rho u_{i}),_{i}=0,$ (53)
$\dfrac{\partial\rho u_{i}}{\partial t}+(\rho u_{i}u_{j})_{,j}=-P_{ij,j}+(\mu(u_{i,j}+u_{j,i}))_{,j},$ (54)
$\dfrac{\partial\rho e}{\partial t}+(\rho eu_{j})_{,j}=-P_{jk}u_{k,j}+\mu(u_{k,j}+u_{j,k})u_{k,j}+(\lambda T_{,j})_{,j},$ (55)
$P_{ij}=\left(p(\rho,T)-\kappa\left(\rho\rho_{,kk}+\frac{1}{2}\rho_{,k}\rho_{,k}-\frac{\rho}{T}\rho_{,k}T_{,k}\right)\right)\delta_{ij}+\kappa\rho_{,i}\rho_{,j}.$ (56)
In addition to density, velocity and viscosity $\rho$, $u_{i}$ and $\mu$, we have internal energy $e$, temperature $T$ and heat conductivity $\lambda$, $\mu$ and $\lambda$ taken constant for simplicity. $P_{ij}$ denotes the pressure tensor, including the Korteweg stress [anderson_AnnRevFluMech1998], and $\kappa$ is the gradient energy coefficient. The function $p(\rho,T)$ is the equation of state.
The equililbrium solution to these equations, corresponding to (35), is found for $u_{i}=0$ and $T=const$, and $P_{ij}=p_{s}(T)\delta_{ij}$, where $p_{s}(T)$ is the liquid-vapor coexistence pressure at temperature $T$. In implicit form this can be written as
$f(\rho,T)+p_{s}(T)=\frac{\kappa}{2}\rho_{x}^{2},$ (57)
where the free energy $f$ is related to the pressure as $p=-f+\rho f_{\rho}$ (indices denoting derivatives).
Analogous to (31)–(38), we now examine the departure from this equilibrium due to interface curvature, and thus derive the Gibbs-Thomson relation for this specific diffuse interface model with a gradient energy. We make a change of variable as before, $r=R+\epsilon x$ where $\epsilon<<1$ indicates that curvature is much larger than the interface width. Introducing perturbed quantities as
$P{ij}$ $=p_{s}(T)\delta{ij}+\epsilon P_{ij}^{1},$ (58)
$\rho$ $=\rho_{0}+\epsilon\rho_{1},$
$u_{i}$ $=0+\epsilon u_{i}^{1},$
$e$ $=e_{0}+\epsilon e_{1},$
$T$ $=T_{s}+\epsilon T_{1}.$
Introducing these in $P_{11}$ and expanding we find
$P_{11}$ $=P_{11}^{0}+\epsilon P_{11}^{1},$ (59)
$P_{11}^{0}$ $=p_{s}(T_{s}),$
$P_{11}^{1}$ $=T_{1}\dfrac{dp_{s}}{dT}+\kappa(\rho_{1}\rho_{0}f_{\rho\rho}+\rho_{0x}\rho_{1x}-\rho_{1}\rho_{0xx}-\rho_{0}\rho_{1xx})+\kappa\frac{1}{R}\rho_{0}\rho_{0x}.$
Multiplying $P_{11}^{1}$ by $\rho_{0x}/\rho_{0}^{2}$ and integrating we can obtain a solvability condition similar to (38). The second term in $P_{11}^{1}$ can be shown to vanish using the relation in (57). The remaining terms are then
$T_{1}\dfrac{dp_{s}}{dT}\int_{v}^{l}\frac{\rho_{0x}}{\rho_{0}^{2}}dx+\kappa\frac{1}{R}\int_{v}^{l}\frac{\rho_{0x}^{2}}{\rho_{0}}dx$ $=0,$ (60)
where
$\int_{v}^{l}\frac{\rho_{0x}}{\rho_{0}^{2}}dx$ $=\frac{1}{\rho_{v}}-\frac{1}{\rho_{l}}=\frac{1}{\tilde{\rho}}$,
$\kappa\int_{v}^{l}\frac{\rho_{0x}^{2}}{\rho_{0}}dx$ $=\int_{v}^{l}\frac{\sqrt{2\kappa(f+p_{s})}}{\rho_{0}}d\rho=\frac{\sigma}{\overline{\rho}}.$
Here we denote the density corresponding to difference in specific volume $\tilde{\rho}$ as $\tilde{\rho}=1/(1/\rho_{v}-1/\rho_{l})$. In the last integral the equilibrium solution in (57) was used, we identified the surface tension as $\sigma=\int_{v}^{l}\sqrt{2\kappa(f+p_{s})}d\rho$, and introduced an averaged density as
$\overline{\rho}=\frac{\int_{v}^{l}\sqrt{2\kappa(f+p_{s})}d\rho}{\int_{v}^{l}\frac{\sqrt{2\kappa(f+p_{s})}}{\rho_{0}}d\rho}.$ (61)
The Clapeyron equation in this context reads
$\dfrac{dp_{s}}{dT}=\frac{L}{T(1/\rho_{v}-1/\rho_{l})}=\frac{L\tilde{\rho}}{T}.$ (62)
Here $L$ denotes the latent heat. The first equation in (60) can then be written as
$T_{1}\dfrac{L}{T}+\frac{1}{R}\frac{\sigma}{\overline{\rho}}=0,$ (63)
i.e. the Gibbs-Thomson relation. We note the explicit expression for the average density in (61), and that this would emphasize the low density of the gas phase.
To proceed, we now derive a scaling for the region in the vicinity of the contact line, with a length scale $\delta$, to be determined. The velocity scale $u$ is the liquid velocity. The lowest order temperature is the equilibrium temperature $T_{s}$, and $T_{1}$ is the perturbation of the temperature. In analogy with (39) we assume that the main balance in the energy equation (54) is between the thermal diffusion $(\lambda T_{,}j)_{,j}$ and the $pdV$ like term $P_{jk}u_{k,j}$, which at first order is $p_{s}u_{k,k}\sim p_{s}(\rho_{l}/\rho_{v})(u/\delta)$. Depending on the fluid properties it is also possible that the thermal diffusion is balanced instead by the convective term on the lefthand side of (55), whch would give slightly different estimates. As in (39) we introduce a nondimensional order unity factor $b$, which is to be calibrated by comparison between the model and full simulations
$p_{s}\frac{\rho_{l}}{\rho_{v}}\frac{bu}{\delta}=\lambda\frac{T_{1}}{\delta^{2}}.$ (64)
In the momentum equation C2 we take the balance between the viscous term and the surface tension, estimating $P_{11}$ as $P_{11}\sim T_{1}(dp_{s}/dT)$ and using (62). As in (40) we introduce a nondimensional order unity parameter $a$
$T_{1}\frac{L\tilde{\rho}}{T_{s}}\frac{1}{\delta}=\mu\frac{au}{\delta^{2}}.$ (65)
Here $1/\tilde{\rho}=(1/\rho_{v}-1/\rho_{l})$.
Equations (63), (64) and (65) are now the counterparts to equations (38), (39) and (40), and are to be solved for $\delta$, $R$ and $T_{1}$. The result of this, corresponding to equations (41), (42) and (38), are
$R=-\frac{1}{a}\frac{\delta}{(\mu u/\sigma)}\frac{\tilde{\rho}}{\overline{\rho}}=-\frac{1}{\sqrt{ab}}\frac{1}{(\mu u/\sigma)}\frac{\tilde{\rho}}{\overline{\rho}}\sqrt{\frac{\rho_{v}}{\rho_{l}}}\sqrt{\frac{\lambda\mu}{\tilde{\rho}p_{s}(L/T_{s})}},$ (66)
$\delta=\sqrt{\frac{a}{b}}\sqrt{\frac{\rho_{v}}{\rho_{l}}}\sqrt{\frac{\lambda\mu}{p_{s}\tilde{\rho}(L/T_{s})}},$ (67)
$T_{1}=u\sqrt{ab}\sqrt{\frac{p_{s}\mu}{\lambda(L/T_{s})}\frac{\rho_{l}}{\rho_{v}\tilde{\rho}}}$. (68)
We can now write down the model for the apparent contact angle, as in (43), which is seen to be very similar:
$\cos(\theta_{\mathrm{a}})=\cos(\theta_{\mathrm{e}})+\dfrac{\delta}{R}=\cos(\theta_{\mathrm{e}})+a\frac{\tilde{\rho}}{\overline{\rho}}\operatorname{Ca},$ (69)
with the capillary number $\operatorname{Ca}=\mu u/\sigma$.